6. A bag contains lemon flavored candies only. Malini takes out one candy without looking into the bag. What is the probability that she takes out

(i) an orange flavored candy?

(ii) a lemon flavored candy?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) We know that the bag only contains lemon-flavored candies. So, The no. of orange flavored candies = 0 ∴ The probability of taking out orange flavored candies = 0/1 = 0 (ii) As there are only lemon flavored candies, P(lemon flavored candies) = 1 (or 100%)